# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ IV.2 Correlation functions for non-equilibrium dynamics and the coherence length
Figure 23: Temperature-dependence growth of the coherence length. Left: coherence length $\xi_{12}$($t_{\mathrm{w}}$) vs Monte Carlo time, see Eq. (24), as computed for size $L=256$, for several temperatures [$T_{\mathrm{g}}=1.1019(29)$], data taken from  Fernández and Martín-Mayor [110]. One Monte Carlo step corresponds to 1 ps in physical time. For $T\geq 1.3$, we reach equilibrium. Right: To a first approximation, see Sec. V.2 for more, in the non-equilibrium regime $\xi$ grows with time as a power law $\xi(t_{\mathrm{w}})\sim t_{\mathrm{w}}^{1/z(T)}$. Joined points stand for $T\leq T_{\mathrm{g}}$. Note the constant value $z(T>T_{\mathrm{g}})\approx 6$ (cyan circles).
@@
$C_{4}(\boldsymbol{r},t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle q_{\boldsymbol{x}}(t_{\mathrm{w}})q_{\boldsymbol{x}+\boldsymbol{r}}(t_{\mathrm{w}})\rangle}.$ (22)
This quantity is self-averaging, hence for a sample of macroscopic size the disorder average can be omitted. The expected behavior at long distance is
@@
where $\xi_{\mathrm{micro}}$ is the size of the coherent regions (the glassy domains), $\theta$ is the so-called replicon exponent [203, 202] and $f(x)$ is a scaling function that decreases faster than exponentially for large values of its argument. Typically [see 43, 44] $\xi_{\mathrm{micro}}$ is taken as the $k=1$ value of
@@
When applying Eq. (24), care must be taken to check for finite-size effects and to deal with the noisy tails of $C_{4}$. This can be done in a self-consistent manner [see 44, 34]. Fig. 23 exhibits the sluggish growth of $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
Figure 24: Dynamic heterogeneities and some non-commuting limits. Top: $C(t,t_{\mathrm{w}})$, see Eq. (25), as a function of $t$ for $t_{\mathrm{w}}=4^{i}$ , $i=3;\dots;16$ (lines, $t_{\mathrm{w}}$ grows from bottom to top). We also plot $\lim_{t_{\mathrm{w}}\to\infty}{C}(t,t_{\mathrm{w}})$ (points) which generates an enveloping curve that ends at $q_{\text{EA}}$ (horizontal line), recall Eq. (26). Bottom: The correlation length $\zeta(C,t_{\mathrm{w}})$, extracted from the correlation function in Eq. (27), as a function of $C^{2}$ (same values of $t_{\mathrm{w}}$ as in the top panel). Note from the top panel that, for fixed $t_{\mathrm{w}}$, $C(t,t_{\mathrm{w}})$ is a monotonically decreasing function of $t$. This length, $\zeta$, indicates the spatial range of the correlations between the changes when the spin configurations at times $t$ and $t_{\mathrm{w}}$ are compared. Figure adapted from Alvarez Baños et al. [13].
Figure 25: The comparison of correlation functions evinces a statics-dynamics equivalence. We plot, with continuous lines, the $C_{2+2}(r,C,t_{\mathrm{w}})$ correlation function of Eq. (27) for several values of $r$, and a fixed $t_{\mathrm{w}}$. $C$ decreases as time $t$ increases, and we show the data for the different values of $t$ as a function of $C^{2}$. This dynamical function coincides with another spatial correlation function $C_{4}(r;L|q)$ (points), computed in equilibrium on a system of size $L$ and conditioned to $q=C$ [see 12, for detailed definitions], if the right equivalence is made between $t_{\mathrm{w}}$ and $L$. In particular, the figure, taken from Alvarez Baños et al. [12], uses a simple statics-dynamics dictionary where non-equilibrium data at time $t_{\mathrm{w}}$ is compared to equilibrium averages for a value of $L$ such that $L=3.7\xi_{12}(t_{\mathrm{w}})$. The main plot shows this comparison for $L=32$, which corresponds to $t_{\mathrm{w}}=2^{31}$, and the inset shows the same for $L=24$, corresponding to $t_{\mathrm{w}}=2^{26}$. The errors in both datasets are comparable and smaller than the point size.
@@
as exemplified in Fig. 24. Sometimes, one changes the variable from $t$ to $C(t,t_{\mathrm{w}})$. This is always possible because $C$ monotonically decreases with $t$ at fixed $t_{\mathrm{w}}$. It is worth stressing that the first limit in Eq. (26) is zero under the assumption of weak ergodicity breaking [89], but a non-zero limit has been recently measured in numerical simulations of spin glasses on Bethe lattices [49].
@@
Finally, one can construct a two-time two-site correlation function that provides spatial information about how the new ordering at later times $t_{\mathrm{w}}+t$ builds by eroding the old ordering pattern at time $t$:
$C_{2+2}(\boldsymbol{r},t,t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle c_{\boldsymbol{x}}(t,t_{\mathrm{w}})c_{\boldsymbol{x}+\boldsymbol{r}}(t,t_{\mathrm{w}})-C^{2}(t,t_{\mathrm{w}})\rangle}\,.$ (27)
@@
where the shorthand $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=\sigma_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma_{\boldsymbol{x}}(t_{\mathrm{w}}+t)$ is used. Note that $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=-1$ if the spin at site $\boldsymbol{x}$ has flipped at time $t_{\mathrm{w}}+t$ as compared to time $t$. We name this a defect. The study of these dynamic heterogeneities was initiated in  Castillo et al. [80, 79], Jaubert et al. [157], Aron et al. [20]. Furthermore, Eq. (27) allows the extraction of the correlation length between defects, $\zeta(t,t_{\mathrm{w}})$, which is shown in the lower panel of Fig. 24 by a direct analogy with Eq. (24) [see 44]. By construction, $\zeta(t,t_{\mathrm{w}})$ is upper-bounded by $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
@@
An important result that we shall use throughout the rest of this review is that these off-equilibrium quantities can be related quantitatively to their equilibrium counterparts. To do so, one needs to establish a statics-dynamics dictionary, which defines a correspondence between a given finite time $t_{\mathrm{w}}$ (off-equilibrium) and a finite size $L$ (equilibrium). One of the first such dictionaries, based on a single length scale, $\xi_{12}$, is illustrated in Fig. 25. As we shall see below in Sec. IV.4, it is necessary to use several length scales to define a comprehensive statics-dynamics equivalence.
@@ IV.3 Connecting aging with the equilibrium picture: the fluctuation-dissipation relations
To better quantify spin glasses’ aging behavior, we describe the connection between the off-equilibrium aging dynamics and the equilibrium picture with many states and a non-trivial $P(q)$.
@@
where the first r.h.s. term enforces the spherical constraint and the noise is white $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2T\delta_{ij}\delta(t-t^{\prime})$.
In addition to the correlation function $C(t,t_{\mathrm{w}})$, Eq. (25), the response function is the key observable:
@@
Self-consistency equations for $C(t,t_{\mathrm{w}})$ and $R(t,t_{\mathrm{w}})$ can be derived through standard functional methods [89]. Space limitations prohibits writing their equations here, but we can briefly discuss their solution. Following a quench from a high temperature ($T>T_{\text{d}}$) to a low temperature ($T<T_{\text{d}}$), [^4] The dynamical critical temperature $T_{\text{d}}$ is defined by the divergence of the equilibrium correlation time scale. the relaxation dynamics do not reach an equilibrium nor a stationary state on any finite time scale. They remain trapped in an aging regime, where functions of two times depend on both times. In the simplest scenario, we have $C(t,t_{\mathrm{w}})=\mathcal{C}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$ and $R(t,t_{\mathrm{w}})=(t+t_{\mathrm{w}})^{-1}\mathcal{R}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$, so that relaxation becomes slower while time increases.
Perhaps the most striking result in the Cugliandolo-Kurchan (CK) solution [89] is that, in the aging regime, response and correlation functions are connected by a generalized fluctuation-dissipation relation (GFDR)
@@
where $\lambda=t_{\mathrm{w}}/(t+t_{\mathrm{w}})$ and $x$ is the Parisi parameter in the equilibrium solution to the model. More specifically, close to the equilibrium free energy $f_{0}$, the number of metastable states grows like $\mathcal{N}(f)\sim\exp[N\beta x(f-f_{0})]$, which in turn generates a non-trivial overlap distribution
@@
Therefore, $x$ is the probability that two replicas at equilibrium are in different states. The connection between the off-equilibrium GFDR and the equilibrium $P(q)$ is foundational to the comprehension of disordered systems from the study of their aging dynamics.
The CK ansatz for the solution to the out-of-equilibrium dynamical equations has been extended by Cugliandolo and Kurchan [90] to the SK model (actually a version of it with soft spins). Even in this more complicated model possessing many diverging time scales [264], the GFDR provides the sought after connection between off-equilibrium dynamics and equilibrium states. Indeed, assuming
@@
Nevertheless, it is fair to remember that the CK solution is based on some crucial hypotheses: weak long-term memory and weak ergodicity breaking. Both are required to decouple the aging dynamics at long times from the initial condition and the fast relaxation at short times. Recently, it has been found that weak ergodicity breaking (which corresponds to the fact that any finite-time configuration is eventually forgotten in the long-time limit) is not always true and must be sometimes replaced by strong ergodicity breaking (meaning that even at very long times the aging dynamics takes place in a restricted space having a positive overlap with the initial condition). Strong ergodicity breaking in off-equilibrium dynamics has been observed in mixed $p$-spin spherical models [117, 118] and in spin-glass models on sparse random graphs [49].
All the above results were obtained for mean-field models. In finite-dimensional models, analytical results are much scarcer. An important one was developed by Franz et al. [122, 123]. These authors considered a generic model with short-range interactions in finite dimensions. The only condition was stochastic stabity, i.e., robustness to small random perturbations (see also Sec. III.1). By studying linear response to the application of long-range perturbations, and assuming that the dynamical averages of one-time quantities (e.g., the energy) converge to their equilibrium values, one can again obtain Eq. (34) relating the equilibrium overlap distribution $P(q)$ to the fluctuation-dissipation ratio
@@
[…]
@@
Usually, the integrated response $\chi(t,t_{\mathrm{w}})$ is measured as the ratio $m/h$ where $m$ is the magnetization at time $t_{\mathrm{w}}+t$ induced by a field $h$ switched on at time $t_{\mathrm{w}}$. To avoid non-linear effects in the response, methods have been invented that measure $\chi(t,t_{\mathrm{w}})$ without switching on the magnetic field [250, 85].
@@
We end this section by presenting the first and most important measurement of the fluctuation-dissipation ratio in experimental spin glasses. Hérisson and Ocio [148] obtained an impressive result by measuring, on the same insulating spin glass at a temperature $T=0.8\,T_{\mathrm{g}}$, both the integrated response $\chi(t,t_{\mathrm{w}})$ and the magnetization autocorrelation $C(t,t_{\mathrm{w}})$. By plotting both quantities parametrically in $t$ for several values of $t_{\mathrm{w}}$ (see Fig. 27), Hérisson and Ocio [148] reported a non-trivial fluctuation-dissipation ratio in experimental spin glasses very similar to the one that was already exhibited in numerical simulations of spin-glass models [202, see, e.g., Fig. 26].
The reader will note, however, that the experimental results in Fig. 27 display a clear time dependence. In other words, the experiment of Hérisson and Ocio [148] was unable to reach the infinite-time limit of Eq. (36). [^5] Herisson and Ocio were well aware of this fact and attempted to extrapolate their results to infinite time. Because very large times mean very large coherence lengths, the sluggish growth of $\xi_{\mathrm{micro}}$ at low temperatures depicted in Fig. 23 suggests that it is highly unlikely that any near-term experiments will be able to reach the limit required in Eq. (36). This raises the question: how should Eq. (36) be used when the limit of large times is unreachable? This is, precisely, the question that we shall address in the next section.

### v2

@@ IV.2 Correlation functions for nonequilibrium dynamics and the coherence length
Figure 24: Temperature-dependent growth of the coherence length. Left: coherence length $\xi_{12}$($t_{\mathrm{w}}$) vs Monte Carlo time, see Eq. (26), as computed for size $L=256$, for several temperatures $T_{\mathrm{g}}=1.1019(29)$], data taken from  Fernández and Martín-Mayor [127]. One Monte Carlo step corresponds to 1 ps in physical time. For $T\geq 1.3$, we reach equilibrium. Right: To a first approximation, see Sec. V.2 for more, in the nonequilibrium regime $\xi$ grows with time as a power law $\xi(t_{\mathrm{w}})\sim t_{\mathrm{w}}^{1/z(T)}$. Joined points stand for $T\leq T_{\mathrm{g}}$. Note the constant value $z(T>T_{\mathrm{g}})\approx 6$ (cyan circles).
@@
$C_{4}(\boldsymbol{r},t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle q_{\boldsymbol{x}}(t_{\mathrm{w}})q_{\boldsymbol{x}+\boldsymbol{r}}(t_{\mathrm{w}})\rangle},$ (24)
where $q_{\boldsymbol{x}}(t_{\mathrm{w}})=\sigma^{(1)}_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma^{(2)}_{\boldsymbol{x}}(t_{\mathrm{w}})$. $C_{4}(\boldsymbol{r},t_{\mathrm{w}})$ is self-averaging, hence for a sample of macroscopic size the disorder average can be omitted. The expected behavior at long distance is
@@
where $\xi_{\mathrm{micro}}$ is the size of the coherent regions (the glassy domains), $\theta$ is the replicon exponent [242, 241], and $f(x)$ is a scaling function that decreases faster than exponentially for large values of its argument. Typically [see 45, 46] $\xi_{\mathrm{micro}}$ is taken as the $k=1$ value of
@@
When applying Eq. (26), care must be taken to check for finite-size effects and to deal with the noisy tails of $C_{4}$. This can be done in a self-consistent manner [see 46, 36]. Fig. 24 exhibits the sluggish growth of $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
Figure 25: Top: $C(t,t_{\mathrm{w}})$, see Eq. (27), as a function of $t$ for $t_{\mathrm{w}}=4^{i}$ , $i=3;\dots;16$ (lines, $t_{\mathrm{w}}$ grows from bottom to top). We also plot $\lim_{t_{\mathrm{w}}\to\infty}{C}(t,t_{\mathrm{w}})$ (points) which generates an enveloping curve that ends at $q_{\text{EA}}$ (horizontal line), recall Eq. (28). Bottom: The correlation length $\zeta(C,t_{\mathrm{w}})$, extracted from the correlation function in Eq. (30), as a function of $C^{2}$ (same values of $t_{\mathrm{w}}$ as in the top panel). Note from the top panel that, for fixed $t_{\mathrm{w}}$, $C(t,t_{\mathrm{w}})$ is a monotonically decreasing function of $t$. This length, $\zeta$, indicates the spatial range of the correlations between the changes when the spin configurations at times $t$ and $t_{\mathrm{w}}$ are compared. Figure adapted from Alvarez Baños et al. [15].
Figure 26: The comparison of correlation functions evinces a statics-dynamics equivalence between equilibrium quantities computed on small systems (dots) and nonequilibrium quantities obtained from large systems (continuous lines). Specifically, the continuous lines are $C_{2+2}(r,C,t_{\mathrm{w}})$, see Eq. (30), for several values of $r$, and a fixed $t_{\mathrm{w}}$. As in Fig. 25—bottom, we change variables from $t$ to the monotonically decreasing $C^{2}(t,t_{\mathrm{w}})=q^{2}$ ($q$ is parameter ranging from 0 to 1), hence data at smaller $q^{2}$ correspond to later times. This dynamical function coincides with the equilibrium spatial correlation function $C_{4}(r;L|q)$ defined in Eq. (31) and plotted with points, if the proper equivalence is made between $t_{\mathrm{w}}$ and $L$. In particular, the figure, taken from Alvarez Baños et al. [14], uses a simple statics-dynamics dictionary where nonequilibrium data at time $t_{\mathrm{w}}$ is compared to equilibrium averages for a value of $L$ such that $L=3.7\xi_{12}(t_{\mathrm{w}})$. The main plot shows this comparison for $L=32$, which corresponds to $t_{\mathrm{w}}=2^{31}$, and the inset shows the same for $L=24$, corresponding to $t_{\mathrm{w}}=2^{26}$. The errors in both datasets are comparable and smaller than the point size.
@@
as exemplified in Fig. 25. It is worth stressing that the first limit in Eq. (28) is zero under the assumption of weak ergodicity breaking [102], but a nonzero limit has been recently measured in numerical simulations of spin glasses on Bethe lattices [51].
A related quantity is the two-time link correlation function
$C_{\text{link}}(t,t_{\mathrm{w}})$ $=$ $\overline{\frac{1}{DN}\sum_{\|\boldsymbol{x}-\boldsymbol{y}\|=1}\,\sigma_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma_{\boldsymbol{y}}(t_{\mathrm{w}})}$
$\quad\times\ \overline{\sigma_{\boldsymbol{x}}(t+t_{\mathrm{w}})\sigma_{\boldsymbol{y}}(t+t_{\mathrm{w}})}$
where $D$ is the space dimension and the sum is restricted to lattice nearest neighbors. $C_{\text{link}}$ conveys information about the density of domain walls in the system. It is very important to change the variable from $t$ to $C(t,t_{\mathrm{w}})$. This is always possible because $C$ monotonically decreases with $t$ at fixed $t_{\mathrm{w}}$. The derivative $\mathrm{d}C_{\text{link}}(t,t_{\mathrm{w}})/\mathrm{d}C^{2}(t,t_{\mathrm{w}})$ is able to distinguish between the RSB theory, which predicts a nonvanishing value for the derivative at long $t_{\mathrm{w}}$, and the droplets view of coarsening where the derivative should vanish [45, 46].
@@
Finally, one can construct a two-time, two-site correlation function that provides spatial information about how the new ordering at later times $t_{\mathrm{w}}+t$ builds by eroding the old ordering pattern at time $t$:
$C_{2+2}(\boldsymbol{r},t,t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle c_{\boldsymbol{x}}(t,t_{\mathrm{w}})c_{\boldsymbol{x}+\boldsymbol{r}}(t,t_{\mathrm{w}})-C^{2}(t,t_{\mathrm{w}})\rangle}\,$, (30)
@@
where the shorthand $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=\sigma_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma_{\boldsymbol{x}}(t_{\mathrm{w}}+t)$ is used. Note that $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=-1$ if the spin at site $\boldsymbol{x}$ has flipped at time $t_{\mathrm{w}}+t$ as compared to time $t_{\mathrm{w}}$. We name this a rearrangement, and Eq. (30) allows the extraction of the correlation length between rearrangements, $\zeta(t,t_{\mathrm{w}})$, which is shown in the lower panel of Fig. 25 by a direct analogy with Eq. (26), see Belletti et al. [46]. By construction, $\zeta(t,t_{\mathrm{w}})$ is upper-bounded by $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
The correlation length $\zeta(t,t_{\mathrm{w}})$, measuring the average size of the regions that rearrange between times $t_{\mathrm{w}}$ and $t_{\mathrm{w}}+t$, is connected to the dynamic heterogeneities that have been measured in spin-glass models with great accuracy [88, 87, 261, 260, 186, 319]. However, a precise connection would require an analysis of local quantities, similar to the one discussed in Sec. VI.1.2 for temperature chaos. Such an analysis is not presently available.
@@
An important result that we shall use throughout the rest of this review is that these off-equilibrium quantities can be related quantitatively to their equilibrium counterparts. To do so, one needs to establish a statics-dynamics dictionary, which defines a correspondence between a given finite time $t_{\mathrm{w}}$ (off-equilibrium) and a finite size $L$ (equilibrium). One of the first such dictionaries, based on a single length scale, $\xi_{12}$, is illustrated in Fig. 26, which employs an equilibrium version of $C_{4}(r,t_{\mathrm{w}})$:
$C_{4}(r;L|q)=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\,\langle q_{\boldsymbol{x}}q_{\boldsymbol{x}+\boldsymbol{r}}\rangle_{q}}\,$, (31)
where the thermal average is restricted only to pairs of replicas of an equilibrated system conditioned to have overlap $q$, recall Eq. (5). Mind that $q$ here is a numerical parameter that we are free to vary between 0 and 1.
As we shall see below in Sec. IV.4, it is necessary to use several length scales to define a comprehensive statics-dynamics equivalence.
@@ IV.3 Connecting aging with the equilibrium picture: the fluctuation-dissipation relations
To better quantify spin glasses’ aging behavior, we describe the connection between the off-equilibrium aging dynamics and the equilibrium picture with many states and a nontrivial $P(q)$.
@@
where the first term on the r.h.s. enforces the spherical constraint, and the noise is white: $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2T\delta_{ij}\delta(t-t^{\prime})$.
In addition to the correlation function $C(t,t_{\mathrm{w}})$, Eq. (27), the response function is a key observable:
@@
Self-consistency equations for $C(t,t_{\mathrm{w}})$ and $R(t,t_{\mathrm{w}})$ can be derived through standard functional methods [102]. Space limitations prohibits writing out their equations here, but we can briefly discuss their solution. Following a quench from a high temperature ($T>T_{\text{d}}$) to a low temperature ($T<T_{\text{d}}$), [^6] The dynamical critical temperature $T_{\text{d}}$ is defined by the divergence of the equilibrium correlation time. the relaxation dynamics do not reach an equilibrium nor a stationary state on any finite time scale. They remain trapped in an aging regime, where functions of two times depend on both times. In the simplest scenario, we have $C(t,t_{\mathrm{w}})=\mathcal{C}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$ and $R(t,t_{\mathrm{w}})=(t+t_{\mathrm{w}})^{-1}\mathcal{R}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$, so that relaxation becomes slower while time increases.
Perhaps the most striking result in the Cugliandolo-Kurchan (CK) solution to the spherical $p$-spin model [102] is that, in the aging regime, response and correlation functions are connected by a generalized fluctuation-dissipation relation (GFDR)
@@
where $\lambda=t_{\mathrm{w}}/(t+t_{\mathrm{w}})$ and $x$ is the Parisi parameter in the equilibrium solution to the model. More specifically, for free energies $f$ close to the equilibrium free energy $f_{0}$, the number of metastable states grows like $\mathcal{N}(f)\sim\exp[N\beta x(f-f_{0})]$, which in turn generates a nontrivial overlap distribution
@@
Therefore, $x$ can be regarded as the probability that two replicas at equilibrium are in different states. The connection between the off-equilibrium GFDR and the equilibrium $P(q)$ is foundational to the comprehension of disordered systems from the study of their aging dynamics.
In models of structural glasses, like the spherical $p$-spin, the quantity $\beta x$ is (mostly) constant in the entire aging regime. For this reason, it is useful to introduce the concept of an effective temperature $T_{\text{eff}}=1/(\beta x)$ in the out-of-equilibrium regime [104]. However, for spin-glass models having multiple and differently diverging time scales in the aging regime, $T_{\text{eff}}$ changes with time, and its interpretation as a temperature requires further assumptions, thoroughly discussed by Montanari and Ricci-Tersenghi [260].
The CK ansatz for the solution to the out-of-equilibrium dynamical equations has been extended by Cugliandolo and Kurchan [103] to the SK model (actually a version of it with soft spins). Even in this more complicated model possessing many diverging time scales [330], the GFDR provides the sought-after connection between off-equilibrium dynamics and equilibrium states. Indeed, assuming
@@
Nevertheless, it is fair to remember that the CK solution is based on some crucial hypotheses: weak long-term memory and weak ergodicity breaking. Both are required to decouple the aging dynamics at long times from the initial condition and the fast relaxation at short times. Recently, it has been found that weak ergodicity breaking —which corresponds to the fact that any finite-time configuration is eventually forgotten in the long-time limit— is not always true and must be sometimes replaced by strong ergodicity breaking, meaning that even at very long times the aging dynamics takes place in a restricted space having a positive overlap with the initial condition. Strong ergodicity breaking in off-equilibrium dynamics has been observed in mixed $p$-spin spherical models [138, 139] and in spin-glass models on sparse random graphs [51].
All the above results were obtained for mean-field models. In finite-dimensional models, analytical results are much scarcer. An important one was developed by Franz et al. [143, 144]. These authors considered a generic model with short-range interactions in finite dimensions. The only condition was stochastic stability, i.e., robustness to small random perturbations (see also Sec. III.1). By studying linear response to the application of long-range perturbations, and assuming that the dynamical averages of one-time quantities (e.g., the energy) converge to their equilibrium values, one can again obtain Eq. (38) relating the equilibrium overlap distribution $P(q)$ to the fluctuation-dissipation ratio
@@
In numerical simulations, measuring the instantaneous response $R(t,t_{\mathrm{w}})$ is not advisable, so integrated responses $\chi(t,t_{\mathrm{w}})=\int_{t_{\mathrm{w}}}^{t}R(t,t^{\prime})dt^{\prime}$ have been used instead [240, 302], relating to the equilibrium $x(q)$ as follows
@@
Usually, the integrated response $\chi(t,t_{\mathrm{w}})$ is measured as the ratio $m/h$ where $m$ is the magnetization at time $t_{\mathrm{w}}+t$ induced by a field $h$ switched on at time $t_{\mathrm{w}}$. To avoid nonlinear effects in the response, methods have been invented that measure $\chi(t,t_{\mathrm{w}})$ without switching on the magnetic field [314, 97].
@@
We end this section by presenting the first and most important measurement of the fluctuation-dissipation ratio in experimental spin glasses. Hérisson and Ocio [173] obtained an impressive result by measuring, on the same insulating spin glass at a temperature $T=0.8\,T_{\mathrm{g}}$, both the integrated response $\chi(t,t_{\mathrm{w}})$ and the magnetization autocorrelation $C(t,t_{\mathrm{w}})$. By plotting both quantities parametrically in $t$ for several values of $t_{\mathrm{w}}$ (see Fig. 28), Hérisson and Ocio [173] reported a nontrivial fluctuation-dissipation ratio in experimental spin glasses very similar to the one that was already exhibited in numerical simulations of spin-glass models [241, see, e.g., Fig. 27].
The reader will note, however, that the experimental results in Fig. 28 display a clear time dependence. In other words, the experiment of Hérisson and Ocio [173] was unable to reach the infinite-time limit of Eq. (40). [^7] Herisson and Ocio were well aware of this fact and attempted to extrapolate their results to infinite time. Because very large times mean very large coherence lengths, the sluggish growth of $\xi_{\mathrm{micro}}$ at low temperatures depicted in Fig. 24 suggests that it is highly unlikely that any near-term experiments will be able to reach the limit required in Eq. (40). This raises the question: how should Eq. (40) be used when the limit of long times is unreachable? This is, precisely, the question that we shall address in the next section.